# Scaling and Load-Balancing Equi-Joins

### v2

@@ 6.1 Splitting the Relations
6.2 When not to Perform Broadcast-Joins?
The second and third joins in Eqn. 5 are executed using Broadcast-Joins. We only discuss the second join, but the logic also applies equally to the third. For the Broadcast-Join assumption to hold, broadcasting the records of the small relation, $\mathcal{S}_{CH}$, over the network has to be faster than splitting the large relation, $\mathcal{R}_{HC}$, among the $n$ executors over the network. From [12], the time to read $\mathcal{S}_{CH}$ and broadcast it follows the Big-$\Theta$ below.
@@ 6.2 When not to Perform Broadcast-Joins?
6.3 The Outer Variants of AM-Join
@@
Table 2: The algorithms for the four sub-joins of Eqn. 5 of the outer variants of AM-Join.
@@
Unlike the state-of-the-art distributed industry-scale algorithms, the different flavors of SkewJoin [21] built on top of Microsoft SCOPE, AM-Join supports all variants of outer-joins without record deduplication or custom partitioning of the relations. Moreover, AM-Join does not require introducing a new outer-join operator variant that understands the semantics of witness tuples, a form of tuples introduced in [21] to extend SkewJoin to outer-joins.
7 Collecting Hot Keys
@@
7.1 How many Hot Keys to Collect?
@@
    $|\mathcal{S}_{CH}|\leq\frac{M}{m_{\mathcal{S}}}$ (6)
@@
    $|\kappa_{\mathcal{R}}|_{max}\leq\frac{|\mathcal{S}_{CH}|}{\left(1+\lambda\right)^{\frac{3}{2}}}$ (7)
@@
$|\kappa_{\mathcal{R}}|_{max}=\min\left(\frac{\min\left(|\mathcal{R}|,\frac{M}{m_{\mathcal{S}}}\right)}{\left(1+\lambda\right)^{\frac{3}{2}}},{{}{\frac{M}{m_{Key}}}}\right)$ (8)
7.2 The Hot Keys Cost
@@
The cost of collecting $|\kappa_{\mathcal{R}}|_{max}$ hot keys using the algorithm in [3] is the cost of scanning the local $\mathcal{R}_{i}$ partitions in parallel, and the cost of merging the hot keys over the network in a tree-like manner. Given $m_{b}$ and $m_{\mathcal{R}}$, the average sizes of a key and a $\mathcal{R}$ record, respectively, the total cost is given by Eqn. 9.
$\Delta_{getHotKeys}=\frac{|\mathcal{R}|\times m_{\mathcal{R}}}{n}+|\kappa_{\mathcal{R}}|_{max}\times m_{b}\times\lambda\times\log(n)$ (9)
@@
We like to contrast collecting the hot keys using a distributed heavy-hitters algorithm to histogramming the relation using a distributed quantiles algorithm [65, 87, 41]. Assuming $\mathcal{R}$ is distributed evenly between the $n$ executors, and given an error rate, $\epsilon$, the local heavy-hitters algorithm uses $\Theta(\frac{1}{\epsilon})$ space [61], while the local quantiles algorithm uses $O(\frac{1}{\epsilon}\log(\epsilon\frac{|\mathcal{R}|}{n}))$ space [34] [^11] This is the theoretical bound on any deterministic single-pass comparison-based quantiles algorithm, even if the algorithms in [65, 87, 41] do not necessarily use it.. Collecting heavy hitters is more precise given the same memory, and incurs less communication when merging the local statistics over the network to compute the global statistics.
8 Evaluation Results
@@
We evaluated the scalability and the handling of skewed data of Tree-Join-Basic, Tree-Join, AM-Join-Basic (i.e., AM-Join using Tree-Join-Basic for joining keys that are hot on both sides of the join), AM-Join, Hash-Join, Broadcast-Join, Full-SkewJoin [21] (the state-of-the-art industry-scale algorithm of the PRPD family), and Fine-Grained partitioning for Skew Data (FGSD-Join) [41] (the state-of-the-art of the key-range-division family). We also evaluated the Spark v3.33.3 .x joins [^12] https://databricks.com/blog/2020/05/29/adaptive-query-execution-speeding-up-spark-sql-at-runtime.html, https://docs.databricks.com/spark/latest/spark-sql/aqe.html., the most advanced open-source shared-nothing algorithm. The two main optimizations of Spark3-Join are (a) dynamically changing the execution plan from Shuffle-Join to Broadcast-Join when one relation can fit in memory, and (b) dynamically combining and splitting data partitions to reach almost equi-size partitions for load balancing. Multicasting is not supported by Spark, the framework we used for evaluation. Evaluating multicasting algorithms (e.g., [69]) is left for future work.
@@
The basic Broadcast-Join algorithm was implemented. This requires either of the input partition to fit in memory. To ensure Broadcast-Join does not run out of memory, the implementation could have been enhanced by pre-splitting one of the to-be relations into multiple sub-relations, such that each sub-relation fits in memory, running a Broadcast-Join between the other relation and each sub-relation, and unioning the results of all the joins. However, the standard implementation of Broadcast-Join was used since these optimizations are neither known to be the default out-of-the-box implementation in major DBMSs, nor typically used by the average DBMS user. For the ExpVar-Join algorithm, the model training time was not accounted for when measuring the runtime of any join. For all adaptive algorithms (Tree-Join and AM-Join and variants, and Full-SkewJoin), and ExpVar-Join the hot keys were collected exactly, which is more time consuming, and does not show the strengths of the algorithms for less skewed data. For AM-Join and variants, the computed hot keys were passed to the employed Tree-Join instead of recomputing them. The sample for FGSD-Join was 10×10$\times$10 × the number of partitions, and the number of partitions was 3×3$\times$3 × to 5×5$\times$5 × the number of executors for all algorithms
@@
The performance of the algorithms on outer-joins was in line with the inner-joins, reported below. However, we could build the outer-join variants of neither Full-SkewJoin, since it depends on Microsoft SCOPE-specific deduplication, nor FGSD-Join, since it was missing in [41]. We compared the outer-join variant of the proposed IB-Join against DER [91] and DDR [27] on Small-Large outer-joins.
@@
The algorithms code and the code for the dataset generation and experiment running was written in Scala v2.11.122.11.122.11.122.11.12 and executed using Spark v2.4.32.4.32.4.32.4.3, except for the Spark v3.33.3 .x experiments that were run on Spark v3.0.23.0.23.0.23.0.2, and the code was compiled using Scala v2.12.122.12.122.12.122.12.12. Spark RDDs were used to store the distributed datasets. Kryo serialization was used to (de)serialize the objects, and all the classes needed were registered with Kryo to minimize the (de)serialization overhead. Custom serialization was used for natural self-joins to avoid (de)serializing the same list of records twice when the two joined lists or sub-lists were one and the same. All the code, including the data generation and the evaluation code. was peer reviewed, and tested for correctness. The code will be available on https://github.com/uber.
@@ 8.1 Estimating $\lambda$
RecordSize $10^{2}$ RecordSize $10^{3}$
NumExecutors $10^{2}$ $7.34$ $7.28$
NumExecutors $10^{3}$ $6.98$ $8.05$
Table 3: The measurement of $\lambda$, the relative runtime of sending data over the network vs. its IO from/to a local disk.
@@
8.2 Experiments on Synthetic Data
@@
Two sizes of records were experimented with, $S=10^{2}$, and $S=10^{3}$ Bytes, representing relatively small and large records. The Zipf-$\alpha$ was varied between $0.0$ and $1.0$, representing uniform to moderately skewed data. Higher values of $\alpha$ are common in many applications [^14] One example is the word frequency in Wikipedia [90].. However, higher values of $\alpha$ are expected to favor algorithms that handle data skew better, namely AM-Join(-Basic) and Tree-Join(-Basic). These four algorithms split the processing of each hot key among multiple executors. Finally, the number of executors used to execute the join was varied from $100$ to $1000$ executors to evaluate how the algorithms utilize more resources to scale to more skewed datasets. All algorithms were allowed $2$ hours to execute each join.
8.2.1 Equi-Join Scalability with Data Skew
Figure 9: The runtime of the equi-join algorithms on $1000$ executors for $D(\alpha,100)$ while varying the Zipf-$\alpha$. Notice the logarithmic scale on the vertical axes.
Figure 10: The runtime of the equi-join algorithms on $1000$ executors for $D(\alpha,1000)$ while varying the Zipf-$\alpha$. Notice the logarithmic scale on the vertical axes
@@ 8.2.1 Equi-Join Scalability with Data Skew
The experiment was repeated with records of size $10^{3}$ Bytes instead of $10^{2}$ Bytes, and the results are reported in Fig. 10151515This is a correction to the prior paper [60]. The size of the records, $m_{\mathcal{R}\mathcal{S}}$, was erroneously stated as $10^{4}$ Bytes in the [60], but they were actually $10^{3}$ Bytes.. For these relatively large records, the runtimes of all the algorithms were comparable until Zipf-$\alpha$ reached $0.3$ with a slight edge for Hash-Join in the Zipf-$\alpha$ $[0.0,0.15]$ range, and FGSD-Join, ExpVar-Join, and Tree-Join-Basic being the slowest algorithms for Zipf-$\alpha$ in the range $[0.0,0.25]$. For the Zipf-$\alpha$ $[0.2,0.3]$ range, none of the algorithms was a consistent winner. AM-Join had the edge for Zipf-$\alpha$ above $0.35$. Full-SkewJoin, Hash-Join, Spark3-Join , FGSD-Join and ExpVar-Join could not finish within the deadline starting at Zipf-$\alpha$ of $0.5$, $0.65$, $0.65$, , $0.7$ and $0.75$, respectively. Even AM-Join-Basic and Tree-Join-Basic did not scale for Zipf-$\alpha$ starting $0.9$ since they could not fit all the relatively large records of the hottest key in the memory of one executor. Only the load-balanced AM-Join and Tree-Join algorithms were able to succeed for all values of Zipf-$\alpha$.
@@
AM-Join and Tree-Join scale almost linearly with the increased skewness, since they are able to distribute the load on the executors fairly evenly. Their basic counterparts still scaled almost linearly but were slower, due to the bottleneck of building the initial index and splitting the hot keys in the first iteration. From Rel. 4, their runtimes are expected to increase with $\ell_{max}$, the frequency of the hottest key, which is impacted by Zipf-$\alpha$. ExpVar-Join scale as well as AM-Join and Tree-Join and sometimes execute faster for all but the skewed relations. The bottleneck of ExpVar-Join is the high memory requirements discussed in § 4.7. Full-SkewJoin and Hash-Join perform relatively well for weakly skewed joins until both the joined relations become mildly skewed (at a Zipf-$\alpha$ of $0.6$ and $0.4$ in Fig. 10 and Fig. 10, respectively). For moderately skewed data, both algorithms failed to produce the results within the deadline, albeit for different reasons. Full-SkewJoin could not load the hot keys and all their records in the memory of all the executors, while the executors of Hash-Join were bottlenecked by the join results of the hottest keys.
@@
8.2.2 Exploring the Equi-Join Parameter Space
@@
8.2.3 Equi-Join Scalability with Computational Resources
@@
Figure 11: The runtime of the equi-join algorithms for $D(0.65,100)$ while varying the number of executors.
Figure 12: The runtime of the equi-join algorithms for a fraction of $D(0.65,100)$ while varying the number of executors. The fraction is the number of executors $/1000$.
@@
There is another phenomenon that is worth highlighting. The difference in the runtime between AM-Join, Tree-Join and their basic counterparts diminished as more executors were used. The reason is the majority of the scalability come from organizing the equi-join execution in multiple stages, which is shared by all the four algorithms.
@@
8.2.4 Speedup of Natural Self-Joins
@@
From Fig. 14, the speedup was roughly $1.67$ across the board regardless of the Zipf-$\alpha$, and the number of executors. There are two observations to be made. First, the speedup decreased as the Zipf-$\alpha$ increased. This can be attributed to the number of Tree-Join iterations which are not impacted by the processing and IO savings discussed in § 4.4. The larger the Zipf-$\alpha$, the more skewed the data, and the larger number of Tree-Join iterations. Scheduling these iterations is a fixed overhead does not enjoy the same savings discussed in § 4.4. The other observation is as the number of executors increased, the savings tended to decrease. This can be attributed to the fixed overhead of communication among the executors as their number increases.
Figure 13: The natural equi-join speedup of Tree-Join for $D(\alpha,100)$ while varying the Zipf-$\alpha$ using various numbers of executors.
Figure 14: The runtime of the Small-Large outer-join algorithms while varying the size of the smaller relation.
8.2.5 Performance on Small-Large Outer-Joins
Fig. 14 shows the runtimes of IB-Join, DER [91] and DDR [27] [^17] For DER and DDR, the Spark implementations provided by the authors of [27] at https://github.com/longcheng11/small-large were used in these experiments. when performing a right-outer join of two relations with records of size $10^{2}$ Bytes, and keys uniformly selected in the range $[1,2\times 10^{5}]$. The larger relation has $10^{8}$ records, while the size of the smaller relation was varied between $10^{5}$ and $10^{6}$ records. The keys in the smaller relation were all even to ensure a selectivity of $50\%$. This is the selectivity that least favors the IB-Join family optimizations in § 5.1.

### v3

@@ 6.1. Splitting the Relations
6.2. When not to Perform Broadcast-Joins?
The second and third joins in Eqn. 5 are executed using Broadcast-Joins. We only discuss the second join, but the logic also applies equally to the third. For the Broadcast-Join assumption to hold, broadcasting the records of the small relation, $\mathcal{S}_{CH}$, over the network has to be faster than splitting the large relation, $\mathcal{R}_{HC}$, among the $n$ executors over the network. From (Bar-Noy and Kipnis, 1994), the time to read $\mathcal{S}_{CH}$ and broadcast it follows the Big-$\Theta$ below.
@@ 6.2. When not to Perform Broadcast-Joins?
6.3. The Outer Variants of AM-Join
Table 2. The algorithms for the four sub-joins of Eqn. 5 of the outer variants of AM-Join.
@@ 6.3. The Outer Variants of AM-Join
Unlike the state-of-the-art distributed industry-scale algorithms, the different flavors of SkewJoin (Bruno et al., 2014) built on top of Microsoft SCOPE, AM-Join supports all variants of outer-joins without record deduplication or custom partitioning of the relations. Moreover, AM-Join does not require introducing a new outer-join operator variant that understands the semantics of witness tuples, a form of tuples introduced in (Bruno et al., 2014) to extend SkewJoin to outer-joins.
7. Collecting Hot Keys
@@
7.1. How many Hot Keys to Collect?
@@
    (6) $|\mathcal{S}_{CH}|\leq\frac{M}{m_{\mathcal{S}}}$
@@
    (7) $|\kappa_{\mathcal{R}}|_{max}\leq\frac{|\mathcal{S}_{CH}|}{\left(1+\lambda\right)^{\frac{3}{2}}}$
@@
(8) $|\kappa_{\mathcal{R}}|_{max}=\min\left(\frac{\min\left(|\mathcal{R}|,\frac{M}{m_{\mathcal{S}}}\right)}{\left(1+\lambda\right)^{\frac{3}{2}}},\frac{M}{m_{Key}}\right)$
7.2. The Hot Keys Cost
@@
The cost of collecting $|\kappa_{\mathcal{R}}|_{max}$ hot keys using the algorithm in (Agarwal et al., 2013) is the cost of scanning the local $\mathcal{R}_{i}$ partitions in parallel, and the cost of merging the hot keys over the network in a tree-like manner. Given $m_{b}$ and $m_{\mathcal{R}}$, the average sizes of a key and a $\mathcal{R}$ record, respectively, the total cost is given by Eqn. 9.
(9) $\Delta_{getHotKeys}=\frac{|\mathcal{R}|\times m_{\mathcal{R}}}{n}+|\kappa_{\mathcal{R}}|_{max}\times m_{b}\times\lambda\times\log(n)$
@@
We like to contrast collecting the hot keys using a distributed heavy-hitters algorithm to histogramming the relation using a distributed quantiles algorithm (Okcan and Riedewald, 2011; Vitorovic et al., 2016; Gavagsaz et al., 2019). Assuming $\mathcal{R}$ is distributed evenly between the $n$ executors, and given an error rate, $\epsilon$, the local heavy-hitters algorithm uses $\Theta(\frac{1}{\epsilon})$ space (Metwally et al., 2005), while the local quantiles algorithm uses $O(\frac{1}{\epsilon}\log(\epsilon\frac{|\mathcal{R}|}{n}))$ space (Cormode and Veselỳ, 2020) [^11] This is the theoretical bound on any deterministic single-pass comparison-based quantiles algorithm, even if the algorithms in (Okcan and Riedewald, 2011; Vitorovic et al., 2016; Gavagsaz et al., 2019) do not necessarily use it.. Collecting heavy hitters is more precise given the same memory, and incurs less communication when merging the local statistics over the network to compute the global statistics.
8. Evaluation Results
@@
We evaluated the scalability and the handling of skewed data of Tree-Join-Basic, Tree-Join, AM-Join-Basic (i.e., AM-Join using Tree-Join-Basic for joining keys that are hot on both sides of the join), AM-Join, Hash-Join, Broadcast-Join, Full-SkewJoin (Bruno et al., 2014) (the state-of-the-art industry-scale algorithm of the PRPD family), and Fine-Grained partitioning for Skew Data (FGSD-Join) (Gavagsaz et al., 2019) (the state-of-the-art of the key-range-division family). We also evaluated the Spark v3.33.3 .x joins [^12] https://databricks.com/blog/2020/05/29/adaptive-query-execution-speeding-up-spark-sql-at-runtime.html, https://docs.databricks.com/spark/latest/spark-sql/aqe.html., the most advanced open-source shared-nothing algorithm. The two main optimizations of Spark3-Join are (a) dynamically changing the execution plan from Shuffle-Join to Broadcast-Join when one relation can fit in memory, and (b) dynamically combining and splitting data partitions to reach almost equi-size partitions for load balancing. Multicasting is not supported by Spark, the framework we used for evaluation. Evaluating multicasting algorithms (e.g., (Polychroniou et al., 2018)) is left for future work.
@@
The basic Broadcast-Join algorithm was implemented. This requires either of the input partitions to fit in memory. To ensure Broadcast-Join does not run out of memory, the implementation could have been enhanced by pre-splitting one of the to-be relations into multiple sub-relations, such that each sub-relation fits in memory, running a Broadcast-Join between the other relation and each sub-relation, and unioning the results of all the joins. However, the standard implementation of Broadcast-Join was used since these optimizations are neither known to be the default out-of-the-box implementation in major DBMSs, nor typically used by the average DBMS user. For the ExpVar-Join algorithm, the model training time was not accounted for when measuring the runtime of any join. For all adaptive algorithms (Tree-Join and AM-Join and variants, and Full-SkewJoin), and ExpVar-Join the hot keys were collected exactly, which is more time consuming, and does not show the strengths of the algorithms for less skewed data. For AM-Join and variants, the computed hot keys were passed to the employed Tree-Join instead of recomputing them. The sample for FGSD-Join was 10×10$\times$10 × the number of partitions, and the number of partitions was 3×3$\times$3 × to 5×5$\times$5 × the number of executors for all algorithms.
@@
The performance of the algorithms on outer-joins was in line with the inner-joins, reported below. However, we could build the outer-join variants of neither Full-SkewJoin, since it depends on Microsoft SCOPE-specific deduplication, nor FGSD-Join, since it was missing in (Gavagsaz et al., 2019). We compared the outer-join variant of the proposed IB-Join against DER (Xu and Kostamaa, 2010) and DDR (Cheng et al., 2017) on Small-Large outer-joins.
@@
The algorithms code and the code for the dataset generation and experiment running was written in Scala v2.11.122.11.122.11.122.11.12 and executed using Spark v2.4.32.4.32.4.32.4.3, except for the Spark v3.33.3 .x experiments that were run on Spark v3.0.23.0.23.0.23.0.2, and the code was compiled using Scala v2.12.122.12.122.12.122.12.12. Spark RDDs were used to store the distributed datasets. Kryo serialization was used to (de)serialize the objects, and all the classes needed were registered with Kryo to minimize the (de)serialization overhead. Custom serialization was used for natural self-joins to avoid (de)serializing the same list of records twice when the two joined lists or sub-lists were one and the same. All the code, including the data generation and the evaluation code, was peer reviewed, and tested for correctness.
@@ 8.1. Estimating $\lambda$
Table 3. The measurement of $\lambda$, the relative runtime of sending data over the network vs. its IO from/to a local disk.
RecordSize = $10^{2}$ RecordSize = $10^{3}$
NumExecutors = $10^{2}$ $7.34$ $7.28$
NumExecutors = $10^{3}$ $6.98$ $8.05$
@@
8.2. Experiments on Synthetic Data
@@
Two sizes of records were experimented with, $S=10^{2}$, and $S=10^{3}$ Bytes, representing relatively small and large records. The Zipf-$\alpha$ was varied between $0.0$ and $1.0$, representing uniform to moderately skewed data. Higher values of $\alpha$ are common in many applications [^14] One example is the word frequency in Wikipedia (Word frequency in Wikipedia ([n. d.]), November 27, 2006).. However, higher values of $\alpha$ are expected to favor algorithms that handle data skew better, namely AM-Join(-Basic) and Tree-Join(-Basic). These four algorithms split the processing of each hot key among multiple executors. Finally, the number of executors used to execute the join was varied from $100$ to $1000$ executors to evaluate how the algorithms utilize more resources to scale to more skewed datasets. All algorithms were allowed $2$ hours to execute each join.
8.2.1. Equi-Join Scalability with Data Skew
Figure 9. The runtime of the equi-join algorithms on $1000$ executors for $D(\alpha,100)$ while varying the Zipf-$\alpha$. Notice the logarithmic scale on the vertical axes.
Figure 10. The runtime of the equi-join algorithms on $1000$ executors for $D(\alpha,1000)$ while varying the Zipf-$\alpha$. Notice the logarithmic scale on the vertical axes
@@ 8.2.1. Equi-Join Scalability with Data Skew
The experiment was repeated with records of size $10^{3}$ Bytes instead of $10^{2}$ Bytes, and the results are reported in Fig. 10151515This is a correction to the prior paper (Metwally, 2022). The size of the records, $m_{\mathcal{R}\mathcal{S}}$, was erroneously stated as $10^{4}$ Bytes in the (Metwally, 2022), but they were actually $10^{3}$ Bytes.. For these relatively large records, the runtimes of all the algorithms were comparable until Zipf-$\alpha$ reached $0.3$ with a slight edge for Hash-Join in the Zipf-$\alpha$ $[0.0,0.15]$ range, and FGSD-Join, ExpVar-Join, and Tree-Join-Basic being the slowest algorithms for Zipf-$\alpha$ in the range $[0.0,0.25]$. For the Zipf-$\alpha$ $[0.2,0.3]$ range, none of the algorithms was a consistent winner. AM-Join had the edge for Zipf-$\alpha$ above $0.35$. Full-SkewJoin, Hash-Join, Spark3-Join, FGSD-Join and ExpVar-Join could not finish within the deadline starting at Zipf-$\alpha$ of $0.5$, $0.65$, $0.65$, $0.7$ and $0.75$, respectively. Even AM-Join-Basic and Tree-Join-Basic did not scale for Zipf-$\alpha$ starting $0.9$ since they could not fit all the relatively large records of the hottest key in the memory of one executor. Only the load-balanced AM-Join and Tree-Join algorithms were able to succeed for all values of Zipf-$\alpha$.
@@
AM-Join and Tree-Join scale almost linearly with the increased skewness, since they are able to distribute the load on the executors fairly evenly. Their basic counterparts still scaled almost linearly but were slower, due to the bottleneck of building the initial index and splitting the hot keys in the first iteration. From Rel. 4, their runtimes are expected to increase with $\ell_{max}$, the frequency of the hottest key, which is impacted by Zipf-$\alpha$. ExpVar-Join scales as well as AM-Join and Tree-Join and sometimes executes faster for all but the skewed relations. The bottleneck of ExpVar-Join is the high memory requirements discussed in § 4.7.Full-SkewJoin and Hash-Join perform relatively well for weakly skewed joins until both the joined relations become mildly skewed (at a Zipf-$\alpha$ of $0.6$ and $0.4$ in Fig. 10 and Fig. 10, respectively). For moderately skewed data, both algorithms failed to produce the results within the deadline, albeit for different reasons. Full-SkewJoin could not load the hot keys and all their records in the memory of all the executors, while the executors of Hash-Join were bottlenecked by the join results of the hottest keys.
@@
8.2.2. Exploring the Equi-Join Parameter Space
@@
8.2.3. Equi-Join Scalability with Computational Resources
@@
Figure 11. The runtime of the equi-join algorithms for $D(0.65,100)$ while varying the number of executors.
Figure 12. The runtime of the equi-join algorithms for a fraction of $D(0.65,100)$ while varying the number of executors. The fraction is the number of executors $/1000$.
@@
There is another phenomenon that is worth highlighting. The difference in the runtime between AM-Join, Tree-Join and their basic counterparts diminished as more executors were used. The reason is the majority of the scalability comes from organizing the equi-join execution in multiple stages, which is shared by all the four algorithms.
@@
8.2.4. Speedup of Natural Self-Joins
@@
From Fig. 14, the speedup was roughly $1.67$ across the board regardless of the Zipf-$\alpha$, and the number of executors. There are two observations to be made. First, the speedup decreased as the Zipf-$\alpha$ increased. This can be attributed to the number of Tree-Join iterations which are not impacted by the processing and IO savings discussed in § 4.4. The larger the Zipf-$\alpha$, the more skewed the data, and the larger number of Tree-Join iterations. Scheduling these iterations is a fixed overhead and does not enjoy the same savings discussed in § 4.4. The other observation is as the number of executors increased, the savings tended to decrease. This can be attributed to the fixed overhead of communication among the executors as their number increases.
Figure 13. The natural equi-join speedup of Tree-Join for $D(\alpha,100)$ while varying the Zipf-$\alpha$ using various numbers of executors.
Figure 14. The runtime of the Small-Large outer-join algorithms while varying the size of the smaller relation.
8.2.5. Performance on Small-Large Outer-Joins
Fig. 14 shows the runtimes of IB-Join, DER (Xu and Kostamaa, 2010) and DDR (Cheng et al., 2017) [^17] For DER and DDR, the Spark implementations provided by the authors of (Cheng et al., 2017) at https://github.com/longcheng11/small-large were used in these experiments. when performing a right-outer join of two relations with records of size $10^{2}$ Bytes, and keys uniformly selected in the range $[1,2\times 10^{5}]$. The larger relation has $10^{8}$ records, while the size of the smaller relation was varied between $10^{5}$ and $10^{6}$ records. The keys in the smaller relation were all even to ensure a selectivity of $50\%$. This is the selectivity that least favors the IB-Join family optimizations in § 5.1.